# Conformally invariant boundary arcs in double dimers

### v1

@@ 3.1.2. Loops and arcs on a Temperleyan approximation
[…]
@@
This is coherent with what we expect of the continuous model: when $z$ approaches the real axis, the number of arcs enclosing $z$ explodes and it is odd or even with probability $1/2$. When $z$ approaches the boundary of the strip, the probability that there is at least one arc enclosing $z$ goes to 00, and so does the probability that there is an odd number of arcs. Besides, the number of arcs above a point is always positive and explodes when this point approaches the real axis.
4. Proof of the main theorem: the folded case.
@@
  • Use the asymptotic expression of $K^{-1}$ in terms of the continuous Green’s function obtained by Kenyon in [15] (see Appendix B) to show that $\mathrm{tr}((K^{-1}S)^{n})$ converges to an $n$-fold holomorphic integral.
@@
The first equality is clear: an arc is either oriented clockwise or counterclockwise. We explain why the second one holds. For $\omega\in\Omega(\mathcal{G}_{\varepsilon})$, the arcs enclosing $z$ are alternating in orientation. To see this, assume that two arcs $A_{1}$ and $A_{2}$ both enclosing $z$ have the same orientation, say clockwise $w_{A_{1}}<w_{A_{2}}<b_{A_{2}}<b_{A_{1}}$, and that there is no other arc enclosing $z$ between them. This cannot happen because there is an odd number of boundary vertices between $w_{A_{1}}$ and $w_{A_{2}}$ so they cannot be matched together. For the same reason, since the leftmost corner on the boundary is black while the rightmost is white by definition of the Temperleyan domain, there is always 00 or $1$ more counterclockwise than clockwise arc which encloses $z$. Hence, $o_{\varepsilon}(z):=l_{\varepsilon}(z)-r_{\varepsilon}(z)=\mathbbm{1}\{n_{\varepsilon}(z)\text{ is odd}\}\in\{0,1\}$, and
@@
To lighten the notation, we write $E_{\varepsilon}:=E_{\varepsilon}(\gamma)$. By definition, the zipper is a set of directed edges: we also say that an undirected edge $wb\in E(\mathcal{G}_{\varepsilon})$ belongs to the zipper and we write $wb\in E_{\varepsilon}$ if $(w,b)\in E_{\varepsilon}$ or $(b,w)\in E_{\varepsilon}$. We also say that a (white or black) vertex $x\in\mathcal{G}_{\varepsilon}$ belongs to the zipper and we write $x\in E_{\varepsilon}$ if there exists $y\in\mathcal{G}_{\varepsilon}$ such that $xy\in E_{\varepsilon}$. If $wb$ is an undirected edge in $\mathcal{G}_{\varepsilon}$, we define its sign by $\text{sgn}(E_{\varepsilon})_{w,b}=\mathbbm{1}\{(w,b)\in E_{\varepsilon}\}-\mathbbm{1}\{(b,w)\in E_{\varepsilon}\}$. We write the vertices of $\mathcal{G}_{\varepsilon}^{r}$ as $(x,i)$ with $x\in\mathcal{G}_{\varepsilon}$, $i\in\{1,2\}$ where $i$ indicates whether $x$ is in the upper or lower half plane. We use the convention that $i=1$ for the vertices on the horizontal axis. We note that $K_{2\alpha}$ is skew-symmetric, $(K_{2\alpha})_{(w,i),(w^{\prime},j)}=(K_{2\alpha})_{(b,i),(b^{\prime},j)}=0$ for any $(w,i),(w^{\prime},j)\in W(\mathcal{G}_{\varepsilon}^{r})$ and $(b,i),(b^{\prime},j)\in\mathcal{G}_{\varepsilon}^{r}$, and that for two vertices $(w,i)\in W(\mathcal{G}_{\varepsilon}^{r}),(b,j)\in B(\mathcal{G}_{\varepsilon}^{r})$,
$(K_{2\alpha})_{(w,i),(b,j)}$ $=\zeta_{w,b}\mathbbm{1}\{w\sim b\}(\mathbbm{1}\{i=j\}+2\alpha\mathbbm{1}\{i=2,j=1\}\text{sgn}(E_{\varepsilon})_{w,b})$
@@
$S_{(w,i),(b,j)}=\mathbbm{1}\{w\sim b,i=2,j=1\}\zeta_{w,b}\text{sgn}(E_{\varepsilon})_{w,b}$
@@
$\tilde{S}_{(w,i),(b,j)}=\mathbbm{1}\{i=2,j=1,w\sim b\}\zeta_{w,b}\text{sgn}(E_{\varepsilon})_{w,b}(-1)^{w\in W_{0}}.$ (21)
@@
$\det(\tilde{K}_{\alpha})=\bigg{(}\prod_{(x,i)\in\mathcal{G}_{\varepsilon}^{r}}\varepsilon(x,i)\bigg{)}^{2}\det(\tilde{K})=\det(\tilde{K}).$
@@
$=\mathbbm{1}\{i=2\}\sum_{b:\{ub\}\in E_{\varepsilon}}\zeta_{u,b}\text{sgn}(E_{\varepsilon})_{u,b}(-1)^{\mathbbm{1}\{u\in W_{0}\}}\tilde{K}^{-1}_{(b,1),(v,j)}.$
@@
$\mathrm{tr}\big{(}(\tilde{S}\tilde{K}^{-1})^{n}\big{)}$ $=2\sum_{w_{0},\dots,w_{n-1}}\sum_{\underset{\{w_{k}b_{k}\}\in E_{\varepsilon}}{b_{0},\dots,b_{n-1}:}}\prod_{k=0}^{n-1}\zeta_{w_{k},b_{k}}\text{sgn}(E_{\varepsilon})_{w_{k},b_{k}}(-1)^{\mathbbm{1}\{w_{k}\in W_{0}\}}\tilde{K}^{-1}_{(b_{k},1),(w_{k+1},2)}$ (26)
$=2\sum_{\underset{e_{k}=\{w_{k}b_{k}\}}{e_{0},\dots,e_{n-1}\in E_{\varepsilon}}}\prod_{k=0}^{n-1}\zeta_{w_{k},b_{k}}\text{sgn}(E_{\varepsilon})_{w_{k},b_{k}}(-1)^{\mathbbm{1}\{w_{k}\in W_{0}\}}\tilde{K}^{-1}_{(b_{k},1),(w_{k+1},2)}.$
@@
We choose a specific path $\gamma$ in $U$ (except for its endpoint $z_{\partial}\in\partial U$) for which the computations are tractable. More precisely, we choose $\gamma$ from $z$ to $z_{\partial}$ polygonal with slope ±$\pm 1$ consisting of a finite number of line segments, and we denote by $E_{\varepsilon}=E_{\varepsilon}(\gamma)$ the associated zipper. For each line segment of $\gamma$ in one of the directions NE, NW, SW, SE, the corresponding portion of the zipper $E_{\varepsilon}(\gamma)$ consists of a zig-zag path of edges, alternately horizontal and vertical. Starting from here, $C=C(U,\gamma)>0$ is a positive constant depending only on $\gamma$ and the open set $U$ that is allowed to change from line to line. The number of edges in the zipper is $|E_{\varepsilon}(\gamma)|=|\gamma|\varepsilon^{-1}+O(1)\leq C\varepsilon^{-1}$.
@@
If we write for all $0\leq k\leq N_{\varepsilon}$, $z_{k}=w_{k}^{0}$ and $dz_{k}=dx_{k}+idy_{k}\in\{\varepsilon(\pm 1\pm i)\}$ the displacement of the corresponding portion of the zipper (NE,NW,SE or SW), $b_{k}^{0}=w_{k}^{0}+(dx_{k}/2,0)$, $w_{k}^{1}=b_{k}^{0}+(0,dy_{k}/2)$, $b_{k}^{1}=w_{k}^{1}+(dx_{k}/2)$,$w_{k+1}^{0}=b_{k}^{1}+(0,dy_{k}/2)$. Observe that $z_{k+1}=w_{k+1}^{0}=z_{k}^{0}+dz_{k}$ except for at most $C(\gamma)$ values of $k$ (corresponding to the last $w_{k}^{0}$ before each change of direction of $\gamma$; recall that $\gamma$ is made of a finite number of straight segments) for which we still have $|z_{k+1}-z_{k}|\leq 2\sqrt{2}\varepsilon$.
@@
$\mathrm{tr}\big{(}(\tilde{S}\tilde{K}^{-1})^{n}\big{)}$ $\approx 2\sum_{\underset{e_{k}=\{w_{k}b_{k}\}}{e_{0},\dots,e_{n-1}\in E_{\varepsilon}\setminus F_{\varepsilon}}}\prod_{k=0}^{n-1}\zeta_{w_{k},b_{k}}\text{sgn}(E_{\varepsilon})_{w_{k},b_{k}}(-1)^{\mathbbm{1}\{w_{k}\in W_{0}\}}\tilde{K}^{-1}_{(b_{k},1),(w_{k+1},2)}$ (28)
$\approx 2\sum_{1\leq k_{1},\dots,k_{n}\leq N_{\varepsilon}}\sum_{\underset{\forall 1\leq i\leq n}{e_{i}=\{w_{i}b_{i}\}\in E_{\varepsilon}(k_{i})}}\prod_{i=0}^{n-1}\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathbbm{1}\{w_{i}\in W_{0}\}}\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)},$
@@
$\left|\mathrm{tr}\big{(}(\tilde{S}\tilde{K}^{-1})^{n}\big{)}-2\sum_{1\leq k_{1},\dots,k_{n}\leq N_{\varepsilon}}\sum_{\underset{\forall 1\leq i\leq n}{e_{i}=w_{i}b_{i}\in E_{\varepsilon}(k_{i})}}\prod_{i=0}^{n-1}\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathbbm{1}\{w_{i}\in W_{0}\}}\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)}\right|$ (29)
@@
$s_{k_{1},\dots,k_{n}}=\sum_{\underset{\forall 1\leq i\leq n}{e_{i}=w_{i}b_{i}\in E_{\varepsilon}(k_{i})}}\prod_{i=0}^{n-1}\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathbbm{1}\{w_{i}\in W_{0}\}}\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)},$
@@
Write $r_{i}=1$ if $w_{i}\in W_{0}$, and $r_{i}=-1$ if $w_{i}\in W_{1}$, and $s_{i}$ similar for $b_{i}$. Then we have, firstly: $(-1)^{\mathbbm{1}\{w_{i}\in W_{0}\}}=-r_{i}$; secondly $\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)}\leq C\varepsilon$ by Equation (27); thirdly Lemma B.3 gives us that there exists $\phi_{\delta}(\varepsilon)\overset{\varepsilon\to 0}{\longrightarrow}0$ such that
@@
indeed, use that $\tfrac{1}{2}(1-r_{i}s_{i})=\mathbbm{1}\{r_{i}\neq s_{i}\}$ and $\tfrac{1}{2}(1+r_{i}s_{i})=\mathbbm{1}\{r_{i}=s_{i}\}$, and the values of $\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}$ in Figure 12. It follows that
$\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathbbm{1}\{w_{i}\in W_{0}\}}=\varepsilon^{-1}\tfrac{1}{2}i[r_{i}dz_{k_{i}}-s_{i}d\overline{z_{k_{i}}}],$
@@
after reindexing $r_{i+1}=r_{i}$. Now similarly to the above, if for some $i=0,\dots,n-1$, we have that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is odd, then the sum over $r_{i}=\pm 1$ gives zero. Hence the only terms that remain have, for all $i=0,\dots,n-1$, that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is even, which is the same as $\mathbbm{1}\{\kappa_{i}\neq\kappa_{i+1}\}=\mathbbm{1}\{\tau=1\}$, which is the same as $\tau_{i}=-(-1)^{\kappa_{i}}(-1)^{\kappa_{i+1}}$. The sum is then independent of $r$, and we obtain
@@
5. Loops and arcs on shifted domains

### v2

@@ 3.1.2. Loops and arcs on a Temperleyan approximation
$c_{2}(z,z_{\partial})=4\Re\left\{\int\int_{\gamma}F_{+}(\overline{z_{1}},z_{0})\overline{F_{+}(\overline{z_{0}},z_{1})}dz_{0}\overline{dz_{1}}-\int\int_{\gamma}F_{-}(\overline{z_{1}},z_{0})F_{-}(\overline{z_{0}},z_{1})dz_{0}dz_{1}\right\}.$
@@
This is coherent with what we expect of the continuous model: when $z$ approaches the real axis, the number of arcs enclosing $z$ explodes and it is odd or even with probability $1/2$. When $z$ approaches the boundary of the strip, the probability that there is at least one arc enclosing $z$ goes to 0, and so does the probability that there is an odd number of arcs. Besides, the number of arcs above a point is always positive and explodes when this point approaches the real axis.
4. Proof of the main theorem: the folded case
@@
  • Use the asymptotic expression of $K^{-1}$ in terms of the continuous Green function obtained by Kenyon in [14] (see Appendix B) to show that $\mathrm{tr}((K^{-1}S)^{n})$ converges to an $n$-fold holomorphic integral.
@@
The first equality is clear: an arc is either oriented clockwise or counterclockwise. We explain why the second one holds. For $\omega\in\Omega(\mathcal{G}_{\varepsilon})$, the arcs enclosing $z$ are alternating in orientation. To see this, assume that two arcs $A_{1}$ and $A_{2}$ both enclosing $z$ have the same orientation, say clockwise $w_{A_{1}}<w_{A_{2}}<b_{A_{2}}<b_{A_{1}}$, and that there is no other arc enclosing $z$ between them. This cannot happen because there is an odd number of boundary vertices between $w_{A_{1}}$ and $w_{A_{2}}$ so they cannot be matched together. For the same reason, since the leftmost corner on the boundary is black while the rightmost is white by definition of the Temperleyan domain, there is always 0 or $1$ more counterclockwise than clockwise arc which encloses $z$. Hence, $o_{\varepsilon}(z):=l_{\varepsilon}(z)-r_{\varepsilon}(z)=\mathds{1}\{n_{\varepsilon}(z)\text{ is odd}\}\in\{0,1\}$, and
@@
To lighten the notation, we write $E_{\varepsilon}:=E_{\varepsilon}(\gamma)$. By definition, the zipper is a set of directed edges: we also say that an undirected edge $wb\in E(\mathcal{G}_{\varepsilon})$ belongs to the zipper and we write $wb\in E_{\varepsilon}$ if $(w,b)\in E_{\varepsilon}$ or $(b,w)\in E_{\varepsilon}$. We also say that a (white or black) vertex $x\in\mathcal{G}_{\varepsilon}$ belongs to the zipper and we write $x\in E_{\varepsilon}$ if there exists $y\in\mathcal{G}_{\varepsilon}$ such that $xy\in E_{\varepsilon}$. If $wb$ is an undirected edge in $\mathcal{G}_{\varepsilon}$, we define its sign by $\text{sgn}(E_{\varepsilon})_{w,b}=\mathds{1}\{(w,b)\in E_{\varepsilon}\}-\mathds{1}\{(b,w)\in E_{\varepsilon}\}$. We write the vertices of $\mathcal{G}_{\varepsilon}^{r}$ as $(x,i)$ with $x\in\mathcal{G}_{\varepsilon}$, $i\in\{1,2\}$ where $i$ indicates whether $x$ is in the upper or lower half plane. We use the convention that $i=1$ for the vertices on the horizontal axis. We note that $K_{2\alpha}$ is skew-symmetric, $(K_{2\alpha})_{(w,i),(w^{\prime},j)}=(K_{2\alpha})_{(b,i),(b^{\prime},j)}=0$ for any $(w,i),(w^{\prime},j)\in W(\mathcal{G}_{\varepsilon}^{r})$ and $(b,i),(b^{\prime},j)\in\mathcal{G}_{\varepsilon}^{r}$, and that for two vertices $(w,i)\in W(\mathcal{G}_{\varepsilon}^{r}),(b,j)\in B(\mathcal{G}_{\varepsilon}^{r})$,
$(K_{2\alpha})_{(w,i),(b,j)}$ $=\zeta_{w,b}\mathds{1}\{w\sim b\}(\mathds{1}\{i=j\}+2\alpha\mathds{1}\{i=2,j=1\}\text{sgn}(E_{\varepsilon})_{w,b})$
@@
$S_{(w,i),(b,j)}=\mathds{1}\{w\sim b,i=2,j=1\}\zeta_{w,b}\text{sgn}(E_{\varepsilon})_{w,b}$
@@
$\tilde{S}_{(w,i),(b,j)}=\mathds{1}\{i=2,j=1,w\sim b\}\zeta_{w,b}\text{sgn}(E_{\varepsilon})_{w,b}(-1)^{w\in W_{0}}.$ (21)
@@
[…]
@@
$=\mathds{1}\{i=2\}\sum_{b:\{ub\}\in E_{\varepsilon}}\zeta_{u,b}\text{sgn}(E_{\varepsilon})_{u,b}(-1)^{\mathds{1}\{u\in W_{0}\}}\tilde{K}^{-1}_{(b,1),(v,j)}.$
@@
$\mathrm{tr}\big((\tilde{S}\tilde{K}^{-1})^{n}\big)$ $=2\sum_{w_{0},\dots,w_{n-1}}\sum_{\underset{\{w_{k}b_{k}\}\in E_{\varepsilon}}{b_{0},\dots,b_{n-1}:}}\prod_{k=0}^{n-1}\zeta_{w_{k},b_{k}}\text{sgn}(E_{\varepsilon})_{w_{k},b_{k}}(-1)^{\mathds{1}\{w_{k}\in W_{0}\}}\tilde{K}^{-1}_{(b_{k},1),(w_{k+1},2)}$ (26)
$=2\sum_{\underset{e_{k}=\{w_{k}b_{k}\}}{e_{0},\dots,e_{n-1}\in E_{\varepsilon}}}\prod_{k=0}^{n-1}\zeta_{w_{k},b_{k}}\text{sgn}(E_{\varepsilon})_{w_{k},b_{k}}(-1)^{\mathds{1}\{w_{k}\in W_{0}\}}\tilde{K}^{-1}_{(b_{k},1),(w_{k+1},2)}.$
@@
We choose a specific path $\gamma$ in $U$ (except for its endpoint $z_{\partial}\in\partial U$) for which the computations are tractable. More precisely, we choose $\gamma$ from $z$ to $z_{\partial}$ polygonal with slope $\pm 1$ consisting of a finite number of line segments, and we denote by $E_{\varepsilon}=E_{\varepsilon}(\gamma)$ the associated zipper. For each line segment of $\gamma$ in one of the directions NE, NW, SW, SE, the corresponding portion of the zipper $E_{\varepsilon}(\gamma)$ consists of a zig-zag path of edges, alternately horizontal and vertical. Starting from here, $C=C(U,\gamma)>0$ is a positive constant depending only on $\gamma$ and the open set $U$ that is allowed to change from line to line. The number of edges in the zipper is $|E_{\varepsilon}(\gamma)|=|\gamma|\varepsilon^{-1}+O(1)\leq C\varepsilon^{-1}$.
@@
If we write for all $0\leq k\leq N_{\varepsilon}$, $z_{k}=w_{k}^{0}$ and $dz_{k}=dx_{k}+idy_{k}\in\{\varepsilon(\pm 1\pm i)\}$ the displacement of the corresponding portion of the zipper (NE,NW,SE or SW), $b_{k}^{0}=w_{k}^{0}+(dx_{k}/2,0)$, $w_{k}^{1}=b_{k}^{0}+(0,dy_{k}/2)$, $b_{k}^{1}=w_{k}^{1}+(dx_{k}/2)$$w_{k+1}^{0}=b_{k}^{1}+(0,dy_{k}/2)$. Observe that $z_{k+1}=w_{k+1}^{0}=z_{k}^{0}+dz_{k}$ except for at most $C(\gamma)$ values of $k$ (corresponding to the last $w_{k}^{0}$ before each change of direction of $\gamma$; recall that $\gamma$ is made of a finite number of straight segments) for which we still have $|z_{k+1}-z_{k}|\leq 2\sqrt{2}\varepsilon$.
@@
$\mathrm{tr}\big((\tilde{S}\tilde{K}^{-1})^{n}\big)$ $\approx 2\sum_{\underset{e_{k}=\{w_{k}b_{k}\}}{e_{0},\dots,e_{n-1}\in E_{\varepsilon}\setminus F_{\varepsilon}}}\prod_{k=0}^{n-1}\zeta_{w_{k},b_{k}}\text{sgn}(E_{\varepsilon})_{w_{k},b_{k}}(-1)^{\mathds{1}\{w_{k}\in W_{0}\}}\tilde{K}^{-1}_{(b_{k},1),(w_{k+1},2)}$ (28)
$\approx 2\sum_{1\leq k_{1},\dots,k_{n}\leq N_{\varepsilon}}\sum_{\underset{\forall 1\leq i\leq n}{e_{i}=\{w_{i}b_{i}\}\in E_{\varepsilon}(k_{i})}}\prod_{i=0}^{n-1}\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathds{1}\{w_{i}\in W_{0}\}}\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)},$
@@
$\left|\mathrm{tr}\big((\tilde{S}\tilde{K}^{-1})^{n}\big)-2\sum_{1\leq k_{1},\dots,k_{n}\leq N_{\varepsilon}}\sum_{\underset{\forall 1\leq i\leq n}{e_{i}=w_{i}b_{i}\in E_{\varepsilon}(k_{i})}}\prod_{i=0}^{n-1}\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathds{1}\{w_{i}\in W_{0}\}}\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)}\right|$ (29)
@@
$s_{k_{1},\dots,k_{n}}=\sum_{\underset{\forall 1\leq i\leq n}{e_{i}=w_{i}b_{i}\in E_{\varepsilon}(k_{i})}}\prod_{i=0}^{n-1}\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathds{1}\{w_{i}\in W_{0}\}}\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)},$
@@
Write $r_{i}=1$ if $w_{i}\in W_{0}$, and $r_{i}=-1$ if $w_{i}\in W_{1}$, and $s_{i}$ similar for $b_{i}$. Then we have, firstly: $(-1)^{\mathds{1}\{w_{i}\in W_{0}\}}=-r_{i}$; secondly $\tilde{K}^{-1}_{(b_{i},1),(w_{i+1},2)}\leq C\varepsilon$ by Equation (27); thirdly Lemma B.3 gives us that there exists $\phi_{\delta}(\varepsilon)\overset{\varepsilon\to 0}{\longrightarrow}0$ such that
@@
indeed, use that $\tfrac{1}{2}(1-r_{i}s_{i})=\mathds{1}\{r_{i}\neq s_{i}\}$ and $\tfrac{1}{2}(1+r_{i}s_{i})=\mathds{1}\{r_{i}=s_{i}\}$, and the values of $\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}$ in Figure 13. It follows that
$\zeta_{w_{i},b_{i}}\text{sgn}(E_{\varepsilon})_{w_{i},b_{i}}(-1)^{\mathds{1}\{w_{i}\in W_{0}\}}=\varepsilon^{-1}\tfrac{1}{2}i[r_{i}dz_{k_{i}}-s_{i}d\overline{z_{k_{i}}}],$
@@
after reindexing $r_{i+1}=r_{i}$. Now similarly to the above, if for some $i=0,\dots,n-1$, we have that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is odd, then the sum over $r_{i}=\pm 1$ gives zero. Hence the only terms that remain have, for all $i=0,\dots,n-1$, that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is even, which is the same as $\mathds{1}\{\kappa_{i}\neq\kappa_{i+1}\}=\mathds{1}\{\tau=1\}$, which is the same as $\tau_{i}=-(-1)^{\kappa_{i}}(-1)^{\kappa_{i+1}}$. The sum is then independent of $r$, and we obtain
@@
5. Proof of the main theorem: the shifted case